OpenStudy (anonymous):

y = x2 + 10x + 11 y = x2 + x − 7

OpenStudy (muzzack):

Geometric figure: Straight Line Slope = -1.636/2.000 = -0.818 x-intercept = -7/9 = -0.77778 y-intercept = -7/11 = -0.63636

OpenStudy (muzzack):

\(\Large\frak\color{red}\ Welcome~To~OS\)

hero (hero):

Looks like a system of equations to me. Let's solve it: y = x^2 + 10x + 11 y = x^2 + x - 7 Subtract the second equation from the first to get: 0 = 9x + 18 -18 = 9x -18/9 = x -2 = x Plug the value of x back in to one of the original equations to find y